# Benchmarking Ultra-Low-Power NPUs

### v2

@@ 3.3. Evaluation Metrics
Latency: Latency is measured using each platform’s internal timer. Notably, all MCUs, bar the MILK-V, are configured to run at 100 MHz. The MILK-V does not support manual frequency scaling, only DVFS. However, latency is inversely proportional to CPU frequency, as described by ( denotes latency, the number of cycles required for a task, and the operating frequency). Accordingly, we normalized MILK-V’s latency to be comparable to performance under uniform frequency conditions. Each model was run for 10 consecutive inferences. We report both the mean latency and standard deviation to account for any run-to-run variability.
Power and Energy: We compute power and energy using the Monsoon High Voltage Power Monitor (msoon, ) at a sampling rate of 50 Hz. The input voltage () is set to 3.3 V, capturing inference duration () and average power (), from which the average energy consumption () is derived. To ensure stable measurements, we analyze only the data recorded after a 1-minute operation period. Power measurements are gathered over 10 repeated inferences, and we report both mean and standard deviation. Fig. 4 details the power profile of YOLOv1 inference on the HX-WE2’s Ethos-U55 NPU.
@@ 3.4. Performance Breakdown
Post-Processing: Any additional operations required to be performed on the CPU. This includes computing softmax outputs for ResidualNet, SimpleNet, and CIFAR10-NAS models. YoloV1 post-processing includes NMS alongside output class softmax. The Autoencoder model does not require post-processing, since it produces direct reconstruction outputs.
@@
The largest for each model is underlined and bolded, while the second largest is in bold.
@@
| YoloV1_small | 2.27±0.004 | 1.76±0.003 | 0.20± 0.005 | 1.83±0.006 | 0.73±0.009 | 0.81±0.008 | 0.01±0.001 | 0.05±0.001 | 0.01±0.001 |
The largest for each model is underlined and bolded, while the second largest is in bold.
@@
| CIFAR10-NAS | 1.11±0.002 | 0.85±0.001 | 0.11±0.002 | 1.09±0.002 | 0.98±0.009 | 1.04±0.008 | 2.75±0.281 |
| ResNet | 1.24±0.003 | 0.85±0.002 | 0.22±0.001 | 2.01±0.002 | 1.08±0.008 | 1.05±0.024 | 9.31±2.567 |
| SimpleNet | 2.30±0.006 | 1.66±0.003 | 0.21±0.007 | 2.13±0.004 | 1.13±0.008 | 1.29±0.010 | 4.13±0.459 |
| Autoenc | 3.94±0.014 | 2.78±0.008 | 6.25±0.203 | 47.06±1.956 | 22.45±0.392 | 12.97±0.232 | 13.74±1.953 |
| YoloV1_small | 2.27±0.004 | 1.76±0.003 | 0.23±0.005 | 1.86±0.007 | 0.91±0.013 | 1.03±0.011 | 5.75±0.770 |
GAP8 | NXP-MCXN947 |
STM32H7A3 | ESP32s3 | |||||||||||||
| Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | |||||
| NAS | 347.67 | 4.96+295.51 | 364.39 | 6.16+295.51 | 358.46 | 534.56 | 569.94 | 371.70 | 127.75 | 551.87 | 127.75 | 538.59 | 423.61 | 93.75 | 674.57 | 268.86 | ||||
| ResNet | 425.38 | 4.98+372.84 | 446.92 | 6.91+372.84 | 258.32 | 372.49 | 471.52 | 381.89 | 127.75 | 618.11 | 127.75 | 694.33 | 456.07 | 70.97 | 694.44 | 268.78 | ||||
| SimpleNet | 214.61 | 5.00+162.55 | 233.04 | 6.87+162.55 | 258.26 | 351.21 | 471.08 | 381.90 | 127.75 | 553.18 | 127.73 | 566.67 | 451.86 | 53.48 | 698.06 | 268.77 | ||||
| Autoenc | 184.15 | 6.46+133.59 | 193.74 | 6.09+133.59 | 143.31 | 196.20 | 261.36 | 381.27 | 125.44 | 336.06 | 125.44 | 336.35 | 203.57 | 21.35 | 445.59 | 271.89 | ||||
| Yolo | 130.43 | 6.93+41.75 | 147.96 | 8.38+41.75 | 43.29 | 159.46 | 287.70 | 410.83 | 152.32 | 263.81 | 152.32 | 319.10 | 119.28 | 167.52 | 355.19 | 268.77 | ||||
@@ 4.1. Power and Efficiency Breakdown
The general-purpose MCUs without dedicated neural hardware – the STM32H7A3ZI and ESP32s3 – demonstrate significantly lower efficiency across all models. This result empirically validates the advantage neural hardware provides for performing on-device inference in constrained environments, with up to 2 orders of magnitude improvement in end-to-end latency in some cases. However, the STM32H7A3ZI’s power consumption during inference (54.91 - 56.11 mW) is comparable to or lower than MAX78000 (C/R) for some models. This is particularly evident for the Autoencoder model, where the STM32H7A3ZI achieves a surprisingly competitive 3.483 – comparable to the best-performing platforms in our suite. This anomaly is likely attributable to the STM32H7A3ZI’s relatively efficient Cortex-M7 core when operating on the Autoencoder’s simple computational structure (0.5455 MMACs). In contrast, the ESP32 consistently exhibits high inference power consumption (129.74 - 157.17 mW) and latency (7.11 – 536.22 ms), despite its advertised support for CPU-accelerated tensor operations. Altogether, while general-purpose MCUs can achieve reasonable efficiency for simple models, they quickly become impractical for more complex NNs.
The MILK-V, our RISC-V SoC, also demonstrates low efficiency across all models, due to its NPU initialization overhead. We observe a different story, however, if initialization overhead is removed from consideration (, for continuous operation). Without initialization, the MILK-V ranks highest for efficiency across almost all benchmark models. Notably, despite a large idle power draw, it achieves blazingly fast inference times (0.17 - 0.61 ms). Thus, for applications where power consumption isn’t a major constraint and workloads don’t require frequent NPU initialization/deinitialization, but low-latency and a compact form factor are key, the MILK-V could prove effective.
@@
The inset graph provides a magnified view of NPU platforms with lower end-to-end latency.
@@ 4.2.2. Memory I/O
Recent work (tinymem, ) has explored optimizing weight loading strategies for such 2D memory layouts to shrink I/O latency when switching models on a single device, including:
virtualizing weight memory within the accelerator to reduce fragmentation,
optimizing dynamic weight allocation to minimize loading/unloading overhead,
and weight preloading, where the next model’s weights are loaded by the idle CPU into unused memory regions before execution.
@@ 4.2.3. Inference
Another unexpected finding is the superior inference latency of the MAX78000 compared to the HX-WE2. The MAX78000 (C), for example, demonstrates an average 2.48x latency improvement of the HX-WE2 (P), despite having significantly lower theoretical compute capacity (30 GOPs vs. 512 GOPs on the HX-WE2). This could be attributed to more optimized weight-stationary dataflow patterns for CNN workloads compared to the Arm Ethos-U55. However, the HX-WE2 still wins in terms of end-to-end latency with much reduced memory I/O latency. The relatively consistent inference times across different models on the HX platforms also suggest its architecture is optimized for larger models than those in our benchmark suite. The MAX78000 demonstrates more variability in inference latency (ranging from 0.14 ms to 4.63 ms), suggesting greater scalability across differing model complexities.
@@ 4.3. Task-Specific Considerations
Operational Modes and Power Profiles The ability to support different operational modes significantly impacts a platform’s suitability for specific applications. The MAX78000 displays high power variation between idle (10.87 - 13.21 mW) and inference (21.13 - 81.67 mW) states; hence, power gating mechanisms could be effectively leveraged in duty-cycled applications.
Further study of the various low-power modes available on our benchmark platforms is needed, including wake-up times, power gating mechanisms, and DVFS. Moreover, dual-CPU platforms with asymmetric co-processing capabilities could improve task distribution between cores — or enable hierarchical task-based wake-up of CPU cores — leading to further power-saving advantages. For instance, MAX78000’s combination of low-power RISC-V and compute-capable Cortex-M4 cores, when used in tandem alongside early-exit strategies for dynamic, low-power inference, could further optimize energy usage during model deployment.
@@ 4.4. Summary of Results
We measured power consumption and latency for various model architectures across commercially-available NPU platforms. We find GOPS isn’t a reliable predictor for estimating end-to-end latency, and memory bandwidth enormously impacts performance.
@@
Excluding initialization overhead (, for applications requiring continuous operation), the MILK-V ranks overall highest in terms efficiency, with its large idle power draw outweighed by fast end-to-end inference latency. Notably, the NXP-MCXN947 ranks highest, both with and without NPU initialization, for the Autoencoder model, delivering over 2× efficiency gains on its nearest competitor.
Key platform differentiators include memory capacity (affecting model complexity), power profiles (, MAX78000 shows significant variation between idle and inference states, beneficial for duty-cycled applications), and precision support (, MAX78000 supports 1-8 bit operations while HX-WE2 supports up to 32-bit).
@@ 4.5. Future Directions
Expanding Operator Support: Currently, most NPU platforms exhibit hugely limited operator support, focusing on CNN-based operators. Future designs should incorporate more expansive operator sets, towards supporting more diverse model architectures, such as transformers.
@@
Enabling On-Device Training: Current NPU platforms exclusively support NN inference, with no support for on-device training. However, model training on-device would enable personalization, continual learning, and adaptation to dynamic distribution shifts, without relying on cloud-based processing – vital for data privacy and remote deployments. Future NPU designs should aim to support quantized on-device training, requiring both memory-efficient training algorithms alongside hardware support for backpropagation.
Standardizing Model Formats: The heterogeneity in supported model formats across our various benchmark platforms is another issue. Vendors should aim to move towards unified model formats to reduce cross-platform compilation and deployment overheads.
@@ 4.6. Practical Recommendations
We offer the following practical recommendations for embedded developers and hardware designers:
For Energy-Efficiency: The MAX78000 largely outperforms other NPU platforms in terms of energy-efficiency (when including NPU initialization), making it particularly well-suited for battery-powered applications. For extended battery life, consider leveraging its ability to dynamically power-gate portions of the system during idle periods.
For Latency-Critical Applications: The HX-WE2 platform offers low-latency with fast NPU initialization, memory I/O, and inference itself, making it best suited for applications requiring responsive model switching, real-time adaptation to changing conditions, or intermittent/duty-cycled operation. The NXP-MCXN947 also achieves low end-to-end inference latency at a significantly lower power budget, making it ideal for power-constrained workloads. Meanwhile, for latency-critical yet space-constrained applications, where power consumption isn’t a major constraint and workloads don’t require frequent NPU initialization/deinitialization, more powerful SoC architectures, like the MILK-V, could also be suitable. Further work could explore the performance of other SoC platforms (luckfox, ; k230, ).
For Large Models: The GAP8’s expansive memory makes it uniquely suitable for deploying larger, more complex models or implementing model-switching approaches where multiple specialized networks are employed based on operating conditions, despite its longer initialization times and inference latency. However, again, if power consumption isn’t a major concern, the MILK-V’s low inference latency and large memory capacity, with SD card support, could also make it a strong alternative.
For Security with Efficiency: Being Arm-based, the NXP-MCXN947 CPU includes Arm’s TrustZone (trustzone, ), enabling hardware isolation between secure and non-secure enclaves. With its competitive end-to-end inference latency (0.96-25.95 ms) and low power draw (22.91–36.69 mW), it may be suitable for security-centric applications without extreme constraints in any one dimension. Future work could explore extending its secure execution environment to integrate NPU acceleration via I/O passthrough, enabling protected yet efficient NN inference.
For Simple Models: For sufficiently simple models, general-purpose MCUs like the STM32H7 can achieve competitive efficiency without dedicated neural acceleration, obviating the need for specialized hardware.

### v3

@@ 3.3. Evaluation Metrics
Measurement Environment: All platforms are measured without concurrent workloads, and all non-essential background processes are disabled where appropriate (, on MILK-V). This minimizes interference during repeated runs. Models are compiled and deployed using each platform’s vendor-provided SDK or toolchain in its default configuration, unless otherwise stated. Exact toolchain versions can be found in our accompanying repository.
Latency: Latency is measured using each platform’s internal timer. Notably, all MCUs, bar the MILK-V, are configured to run at 100 MHz. The MILK-V lacks support for fixed frequency scaling, only DVFS. However, latency is inversely proportional to CPU frequency, as described by (where denotes latency, the number of cycles required for a task, and the operating frequency). Accordingly, we normalized MILK-V’s latency to be comparable to performance under uniform frequency conditions.
Each model was run for 10 consecutive inferences. We report mean latency and standard deviation to account for run-to-run variability. We observe higher variance on the MILK-V SoC, primarily from CMA activity and residual background services. To reduce noise, we increased the number of runs for latency measurement from 10 to 100; while some variability remains, this averaging ensures stable estimates. Other platforms show limited variance under bare-metal execution, so 10 runs are sufficient.
Power and Energy: We compute power and energy using a Monsoon High Voltage Power Monitor (msoon, )with PowerTool v5.0.0.10 software at a sampling rate of 50 kHz. The input voltage () is fixed at 3.3 V. We capture inference duration () and average power () to compute mean energy consumption (). To ensure steady-state measurements, readings are taken only after a 60 s warm-up period. Measurements are repeated for 10 inference runs, and mean ± standard deviation are reported. Fig. 4 shows an example power profile for YOLOv1 inference on the HX-WE2’s Ethos-U55 NPU.
@@ 3.4. Performance Breakdown
Post-Processing: Any operations required to be performed on the CPU. This includes computing softmax outputs for ResidualNet, SimpleNet, and CIFAR10-NAS models. YoloV1 post-processing includes NMS with output class softmax.
@@
The largest for each model is underlined and bolded, while the second largest is bold.
@@
| YoloV1 | 2.27±0.004 | 1.76±0.003 | 0.20±0.005 | 1.83±0.006 | 0.73±0.009 | 0.81±0.008 | 0.01±0.001 | 0.05±0.001 | 0.01±0.001 |
The largest for each model is underlined and bolded, while the second largest is bold.
@@
| CIFAR10-NAS | 1.11±0.002 | 0.85±0.001 | 0.11±0.002 | 1.09±0.002 | 0.98±0.009 | 1.04±0.008 | 2.80±0.077 |
| ResNet | 1.24±0.003 | 0.85±0.002 | 0.22±0.001 | 2.01±0.002 | 1.08±0.008 | 1.05±0.024 | 4.51±0.469 |
| SimpleNet | 2.30±0.006 | 1.66±0.003 | 0.21±0.007 | 2.13±0.004 | 1.13±0.008 | 1.29±0.010 | 4.17±0.195 |
| Autoenc | 3.94±0.014 | 2.78±0.008 | 6.25±0.203 | 47.06±1.956 | 22.45±0.392 | 12.97±0.232 | 13.29±0.883 |
| YoloV1 | 2.27±0.004 | 1.76±0.003 | 0.23±0.005 | 1.86±0.007 | 0.91±0.013 | 1.03±0.011 | 5.75±0.274 |
@@ 4.1. Power and Efficiency Breakdown
The general-purpose MCUs without dedicated neural hardware – the STM32H7A3ZI and ESP32s3 – demonstrate significantly lower efficiency across all models. This result empirically validates the advantage neural hardware provides for performing on-device inference in constrained environments, with up to 2 orders of magnitude improvement in end-to-end latency in some cases. However, the STM32H7A3ZI’s power consumption during inference (54.91 - 56.11 mW) is comparable to or lower than MAX78000 (C/R) for some models. This is particularly evident for the AI8XAutoEnc model, where the STM32H7A3ZI achieves a surprisingly competitive 3.483 – comparable to the best-performing platforms in our suite. This is consistent with its architectural characteristics; it contains only two lightweight 1D convolutional layers followed by a predominantly fully-connected decoder, meaning its workload is compute-light and exhibits minimal parallelism for NPUs to exploit. As such, optimized scalar or SIMD execution on a high-performance MCU core (, STM32H7’s Cortex-M7) can achieve efficiency close to specialized hardware. In contrast, the ESP32 consistently exhibits high inference power consumption (129.74 - 157.17 mW) and latency (7.11 – 536.22 ms), despite its advertised support for CPU-accelerated tensor operations. Altogether, while general-purpose MCUs can achieve reasonable efficiency for simple models, they quickly become impractical for more complex NNs.
The MILK-V, our RISC-V SoC, also demonstrates low efficiency across all models, due to its NPU initialization overhead. We observe a different story, however, if initialization overhead is removed from consideration (, for continuous operation). Without initialization, the MILK-V ranks highest for efficiency across almost all benchmark models. Notably, despite a large idle power draw, it achieves blazingly fast inference times (0.17 - 0.61 ms).
@@ 4.2.2. Memory I/O
| MAX78k (C) | MAX78k (R) | GAP8 | NXP-MCXN947 | HX-WE2 (S) | HX-WE2 (P) | STM32H7A3 | ESP32s3 | |||||||||
| Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | Flash | RAM | |
| NAS | 347.67 | 4.96+295.51 | 364.39 | 6.16+295.51 | 358.46 | 534.56 | 569.94 | 371.70 | 127.75 | 551.87 | 127.75 | 538.59 | 423.61 | 93.75 | 674.57 | 268.86 |
| ResNet | 425.38 | 4.98+372.84 | 446.92 | 6.91+372.84 | 258.32 | 372.49 | 471.52 | 381.89 | 127.75 | 618.11 | 127.75 | 694.33 | 456.07 | 70.97 | 694.44 | 268.78 |
| SimpleNet | 214.61 | 5.00+162.55 | 233.04 | 6.87+162.55 | 258.26 | 351.21 | 471.08 | 381.90 | 127.75 | 553.18 | 127.73 | 566.67 | 451.86 | 53.48 | 698.06 | 268.77 |
| Autoenc | 184.15 | 6.46+133.59 | 193.74 | 6.09+133.59 | 143.31 | 196.20 | 261.36 | 381.27 | 125.44 | 336.06 | 125.44 | 336.35 | 203.57 | 21.35 | 445.59 | 271.89 |
| YoloV1 | 130.43 | 6.93+41.75 | 147.96 | 8.38+41.75 | 43.29 | 159.46 | 287.70 | 410.83 | 152.32 | 263.81 | 152.32 | 319.10 | 119.28 | 167.52 | 355.19 | 268.77 |
@@
Recent work (tinymem, ) has explored optimizing weight loading strategies for such 2D memory layouts to shrink I/O latency when switching models on a single device, including virtualizing weight memory within the accelerator to reduce fragmentation, optimizing dynamic weight allocation to minimize loading/unloading overhead,and weight preloading, where the next model’s weights are loaded by the idle CPU into unused memory regions before execution.
@@ 4.2.3. Inference
Another unexpected finding is the superior inference latency of the MAX78000 compared to the HX-WE2. The MAX78000 (C), for example, demonstrates an average 2.48× latency improvement over the HX-WE2 (P), despite the HX-WE2’s much higher advertised peak compute capacity at its maximum rated frequency (512 GOP/s at 1 GHz vs. 30 GOP/s for the MAX78000). In our benchmark, however, both platforms are operated at 100 MHz, so these peak figures are not directly comparable. The observed advantage may be attributed to more optimized weight-stationary dataflow patterns for CNN workloads compared to the Arm Ethos-U55. However, the HX-WE2 still wins in terms of end-to-end latency with much reduced memory I/O latency. The relatively consistent inference times across different models on the HX platforms also suggest its architecture is optimized for larger models than those in our benchmark suite. The MAX78000 demonstrates more variability in inference latency (ranging from 0.14 ms to 4.63 ms), suggesting greater scalability across differing model complexities.
@@
Architectural factors help explain scaling patterns. Models like CIFAR10-NAS, generated via OFA NAS, make heavy use of frequent 1×1 convolutions and irregular channel scaling — operations that can map efficiently to NPUs with optimized channel-mixing kernels, but which can be memory-bound if channel-width transitions cause repeated buffer reallocations. ResidualNet and SimpleNet share identical convolutional footprints and late-stage expansion to 512 channels, but ResidualNet’s skip connections extend activation lifetimes and reduce buffer reuse, often increasing both memory pressure and execution time on platforms without aggressive activation scheduling.
@@ 4.3. Task-Specific Considerations
Peak activation RAM and lifetime pressure metrics (Table 2) highlight why certain models stress specific platforms more than others. For example, CIFAR10-NAS’s high lifetime pressure (18.39 MBsteps) results in staggered channel scaling and interleaved pooling, which hold intermediate activations live for long durations. This amplifies SRAM demand and can degrade throughput on platforms with smaller or rigidly partitioned on-chip memories.
Operational Modes and Power Profiles The ability to support different operational modes significantly impacts a platform’s suitability for specific applications. The MAX78000 displays high power variation between idle (10.87 - 13.21 mW) and inference (21.13 - 81.67 mW) states; hence, power gating – shutting down unused compute domains or memory banks – could extend its battery life in duty-cycled or event-driven inference scenarios. Moreover, dual-CPU platforms with asymmetric co-processing capabilities could improve task distribution between cores – or enable hierarchical wake-up of CPU cores – leading to power-saving advantages. For instance, MAX78000’s combination of RISC-V and Cortex-M4 cores, when used in tandem alongside early-exit strategies for dynamic, low-power inference, could further optimize energy usage during model deployment.
Our measurements suggest that platforms with large differences between idle and active power (e.g., MAX78000) stand to benefit more from aggressive gating than those with high static draw (e.g., HX-WE2).
@@ 4.4. Summary of Results
We measured power consumption and latency for various model architectures across commercially-available NPU platforms. We find GOPS isn’t a reliable predictor for estimating end-to-end latency, and memory bandwidth enormously impacts performance. The architectural footprint of each model modulates its real-world performance. Memory-bound designs with high activation lifetime (, ResidualNet, CIFAR10-NAS) tend to underutilize available GOPS, whereas depth-heavy but narrow networks (, YOLOv1) appear to approach peak throughput on accelerators with low memory I/O cost. Models with limited convolutions see diminishing returns from NPU offload, explaining cases where MCUs perform competitively.
@@
Excluding initialization overhead (, for applications requiring continuous operation), the MILK-V ranks overall highest in terms efficiency, with its large idle power draw outweighed by fast end-to-end inference latency.
@@ 4.5. Future Directions
Expanding Operator Support: Currently, most NPU platforms exhibit very limited operator support, heavily optimized for convolutional layers and a small set of accompanying functions such as pooling, elementwise activation, and basic addition. This specialization allows aggressive hardware acceleration for CNN-based workloads, but it comes at the cost of excluding operators central to other model architectures. For example, transformer-based architectures require efficient support for dense matrix multiplication (with high-rank tensors), softmax, normalization layers, and attention-specific dataflows – none of which are implemented in NPU instruction sets. In principle, unsupported layers could be offloaded to the CPU, but such heterogeneous execution requires a tightly integrated CPU/NPU software stack capable of minimizing data transfer and synchronization overheads. Notably, this kind of fine-grained CPU offloading is not supported on most of the evaluated platforms; as a result, our benchmarking is necessarily constrained to CNN variants that map well onto the hardware primitives.
@@
Enabling On-Device Training:
Current NPU architectures are optimized exclusively for inference and lack any hardware or software provisions for on-device training. This would enable privacy-preserving domain adaptation in
bandwidth- or connectivity-constrained environments. This will require both architectural support for backpropagation alongside memory-efficient training algorithms capable of operating within the stringent compute, storage, and energy constraints inherent to NPU-class hardware.
Standardizing Model Formats: The heterogeneity in supported model formats across our various benchmark platforms is another issue. Vendors should aim to move towards unified model formats to reduce cross-platform overheads.
@@ 4.6. Practical Recommendations
We offer the following practical recommendations to embedded developers and hardware designers:
For Energy-Efficiency: The MAX78000 largely outperforms other NPU platforms in terms of energy-efficiency (when including NPU initialization overhead), making it particularly well-suited for low- and battery-powered applications. For extended battery life, consider leveraging its ability to power-gate portions of the system during idle periods.
For Latency-Critical Applications: The HX-WE2 platform offers low-latency with rapid NPU initialization, memory I/O, and inference itself, making it best suited for applications requiring responsive model switching, real-time adaptation to changing conditions, or intermittent/duty-cycled operation. The NXP-MCXN947 also achieves relatively low end-to-end inference latency, at a significantly lower power budget, making it ideal for power-constrained workloads. Meanwhile, for latency-critical but space-constrained applications – where power consumption is less of a concern and workloads avoid frequent NPU initialization – developers should explore SoC architectures. Further evaluation is needed for NPU-equipped SoC platforms (luckfox, ; k230, ).
For Large Models: The GAP8’s expansive memory makes it uniquely suitable for deploying larger models, or model-switching approaches where multiple specialized NNs are employed based on operating conditions (despite its longer initialization times and inference latency). However, again, if power consumption isn’t a major concern, SoC-type platforms, with their low inference latency and larger memory capacity, could be a strong alternative.
For Conv-Lite Models: For models with limited convolutional layers, general-purpose MCUs can achieve competitive efficiency without dedicated neural acceleration, potentially eliminating the need for specialized hardware.